OpenStudy (anonymous):

when 36 is subtracted from the square of a number, the result is five times the number. what is the positive solution ?

OpenStudy (anonymous):

think about it. Call your number x and express the sentence in the form of an equation.

OpenStudy (anonymous):

post what you got here.

OpenStudy (anonymous):

ive got nothing.

OpenStudy (anonymous):

would it be like , x-36=5x ?

OpenStudy (anonymous):

you are almost there. are you sure the LHS is x-36?

OpenStudy (anonymous):

x squared?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

then bring over the 5x to the LHS and you have got yourself a quadratic equation. solve for x. The positive root is your answer.

OpenStudy (anonymous):

when i bring the 5x over would the equation be : xsquared-5x-36?

myininaya (myininaya):

yes but dont forget the = 0

myininaya (myininaya):

and thats not a face I tried to type

OpenStudy (anonymous):

lol okay

OpenStudy (anonymous):

yes, this is an equation. you always have a LHS and RHS for an equation.

myininaya (myininaya):

you can see if it is factorable or you can use the quadratic formula

OpenStudy (radar):

It is factorable\[x ^{2}-5x-36=0\]\[(x-9)(x+4)=0\] Take it from there and get the solution(s)